OpenStudy (anonymous):

The pair of polygons is similar. Find the value of x. one triangle is 75 in and 30 in the other is 12 in and x

OpenStudy (anonymous):

i dislike this math unit

OpenStudy (anonymous):

i feel like the answer is 37.5 in but tht seems to big

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Well, that's a good instinct... that doesn't sound right to me either. Here's how I look at it. 75 inches and 30 inches ----->>>>> 12 inches and x inches So, "x" must be less than half of 12, since 30 is less than half of 75. That will at least give you a reasonable idea of whether your answer makes any sense.

OpenStudy (anonymous):

Again, it's a ratio: BIG : LITTLE 75 : 12 30 : ???

OpenStudy (anonymous):

However, in this case, it might be easier to set up the ratio like this: Triangle 1 / Triangle 2 75 / 30 12 / x

OpenStudy (anonymous):

soo uh 30 x 12 = 360

OpenStudy (anonymous):

yes... you can cross multiply like that: 12 * 30 = 75 * x

OpenStudy (anonymous):

yup so 75 x 5 = 375 which is why i thought the other answer

OpenStudy (anonymous):

12 * 30 = 75 * x 360 = 75 * x Then divide both sides by 75: 360 / 75 = x

OpenStudy (anonymous):

4.8 and 1

OpenStudy (anonymous):

x = 4.8) So, to check, find both ratios... they should be equal.: 75/30 = 2.5 and 12 / 4.8 = 2.5 So if x = 4.8, the ratios work out the same.

OpenStudy (anonymous):

this question is mutiple choice and these r the possible answers 187.5 12 in 30 in 37.5 in

OpenStudy (anonymous):

yikes. Do you have a diagram?

OpenStudy (anonymous):

ya lol ya theres a pic

OpenStudy (anonymous):

I apparently mis-imagined what the picture was. My math is right but not for this problem :) If you can post the pic, let's try it again :)

OpenStudy (anonymous):



OpenStudy (anonymous):

looks like tht

OpenStudy (anonymous):

Are the 75, 30, 12, and x labeled on your picture?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you post the pic again (you can use the copy previous picture button) and add the labels?

OpenStudy (anonymous):